# SageAttention2: Efficient Attention with Thorough Outlier Smoothing and Per-thread INT4 Quantization

### v4

@@ Abstract
Although quantization for linear layers has been widely used, its application to accelerate the attention process remains limited. To further enhance the efficiency of attention computation compared to SageAttention while maintaining precision, we propose SageAttention2, which utilizes significantly faster 4-bit matrix multiplication (Matmul) alongside additional precision-enhancing techniques. First, we propose to quantize matrices $(Q,K)$ to INT4 in a hardware-friendly thread-level granularity and quantize matrices $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$, enhancing the accuracy of INT4 $QK^{\top}$. Third, we propose a two-level accumulation strategy for $\widetilde{P}V$ to enhance the accuracy of FP8 $\widetilde{P}V$. The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 4.5x on RTX4090, respectively. Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, while delivering much higher accuracy. Comprehensive experiments confirm that our approach incurs negligible end-to-end metrics loss across diverse models—including those for language, image, and video generation. The code is available at https://github.com/thu-ml/SageAttention
@@ 1 Introduction
Due to the quadratic time complexity of attention, its efficient implementation becomes crucial as sequence lengths increase in real-world applications (Jiang et al., 2024). Several strategies have been developed to mitigate computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2021; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., 2022; Xiao et al., 2024b, a; Chen et al., 2024; Jiang et al., 2024; Venkataramanan et al., 2024; Gao et al., 2024; Fu et al., 2024) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks. The widely used attention methods optimize attention by exploiting hardware capacities, such as FlashAttention (Dao et al., 2022), FlashAttention2 (Dao, 2024), FlashAttention3 (Shah et al., 2024), xformers (Lefaudeux et al., 2022), and SageAttention (Zhang et al., 2025). These works do not omit computations across the entire sequence and achieve impressive speed and accuracy performance across various tasks.
@@
Motivation. For the two matrix multiplication (Matmul) operations in attention: $QK^{\top}$ and $\widetilde{P}V$, SageAttention accelerates them by quantizing the $QK^{\top}$ to INT8 and uses FP16 Matmul with FP16 accumulators for $\widetilde{P}V$. Moreover, to keep the accuracy of attention, SageAttention proposes smoothing $K$ by eliminating its channel-wise outliers. SageAttention achieves 2 $\times$ speedup than FlashAttention2 and is the first quantized attention that incurs negligible end-to-end metrics loss across language, image, and video generation models. However, SageAttention has two weaknesses. (W1) INT8 Matmul achieves only half the speed of INT4. (W2) FP16 Matmul with FP16 accumulators provides a speedup only on RTX 4090 and RTX 3090 GPUs. To leverage the faster INT4 tensor cores for $QK^{\top}$ and using a method that can accelerate $\widetilde{P}V$ on a broader range of GPUs, we propose to quantize $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
@@
Challenges. Quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8 presents significant challenges. For example, when only per-tensor quantizing $Q,K$ to INT4, the text-to-video model CogvideoX will generate a completely blurry video, and Llama3 only achieves a random-guessing-level accuracy of 25% on MMLU. After investigating deeply, we identified three primary challenges: (C1) The numerical range of INT4 is quite limited compared to FP16 or INT8, leading to significant quantization errors when $Q$ and $K$ have some abnormal values. (C2) We discover that the FP32 accumulator designed for FP8 matrix multiplication in the tensor core (mma.f32.f8.f8.f32) is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. This will lead to an accuracy loss of $\widetilde{P}V$.
@@
Our approach. To address (C1), we first discover that the per-block quantization of $Q,K$ in SageAttention does not achieve sufficient accuracy for INT4 quantization. Simultaneously, to avoid the extra latency caused by per-token dequantization, where each GPU thread corresponds to multiple quantization scales, we propose an adequately precise quantization method that incurs no additional dequantization overhead. Specifically, we introduce a per-thread quantization approach based on the mapping between the GPU threads and memory layout of matrices as dictated by the PTX mma instructions. This method groups tokens corresponding to the same thread for quantization and dequantization, ensuring that each thread is associated with a single quantization scale. This approach achieves much better accuracy performance than per-block quantization with no additional dequantization overhead. Second, for the significant channel-wise outliers in matrices $Q$ and $K$, we adopt smoothing $K$ in SageAttention and further propose an effective method to remove these outliers in $Q$. Specifically, we propose subtracting the average value of the channel dimension of $Q$, referred to as $\overrightarrow{Q}_{m}$. Then, we add $\overrightarrow{Q}_{m}K$ after the $QK^{\top}$ Matmul to ensure the correctness of the attention computation. To address (C2), the accuracy loss from the 22-bit accumulator for FP8 Matmul of $\widetilde{P}V$, we propose a two-level accumulation strategy that uses an FP32 buffer to accumulate the values from the 22-bit accumulator after each block Matmul of $\widetilde{P}V$. This method confines the errors to the block range. Additionally, we design an optional technique to enhance the accuracy of the 22-bit accumulator. Specifically, we could smooth $V$ by subtracting the average value of its channel dimension and adding the subtracted item to the attention output to maintain the correctness.
@@
Performance. Importantly, we offer a high-performance implementation of SageAttention2 on RTX4090 and L20 GPUs. This implementation achieves a peak performance of 481 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3x and 4.5x, respectively. Note that FlashAttention3 is tailored to and can only be used with the Nvidia Hopper architecture. Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, while delivering much better accuracy. For example, on CogVideoX-1.5 (Yang et al., 2025), Mochi (Team, 2024) and HunyuanVideo (Kong et al., 2024), our method does not compromise end-to-end accuracy, whereas videos generated using FlashAttention3(fp8) often suffer noticable degradation, as visualized in Fig. 8. We extensively evaluate the end-to-end metrics of state-of-the-art text, image, and video generation models using SageAttention2. SageAttention2 can accelerate models in a plug-and-play way with negligible loss in end-to-end metrics.
@@ 2.1 FlashAttention
The attention computation can be formulated as: $S=QK^{\top}/\sqrt{d},~{}P=\sigma(S),~{}O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$. The matrices $Q$, $K$, and $V$ each has dimensionality $N\times d$, and $S$, $P$ are $N\times N$. $d$ is typically small, e.g., 64 or 128, and $N$ can be thousands or millions. The time complexity of attention is $O(N^{2})$, primarily due to two matrix multiplications ($QK^{\top}$ and $PV$), both with complexities of $O(N^{2}d)$. FlashAttention (Dao, 2024) is a GPU-friendly attention implementation, which tiles $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\}_{i=1}^{n_{q}},\{K_{i}\}_{i=q}^{n_{k}},\{V_{i}\}_{i=1}^{n_{v}}$ with block sizes of $b_{q}$, $b_{k}$, $b_{v}$ tokens, respectively, where $n_{q},n_{k},n_{v}$ are the number of tiles, and $b_{k}$ = $b_{v}$. FlashAttention computes the output matrix $O$ in parallel in tiles. Each streaming multiprocessor (SM) computes a block $O_{i}$ (corresponds to a $Q_{i}$) by iteratively loads $K_{j},V_{j}$ for each $j$, and update the output with online softmax (Milakov & Gimelshein, 2018):
@@
where $m_{ij}$ and $l_{ij}$ are $b_{q}$-dimenalional vectors, initialized with $\infty$- ∞ and 00 respectively. $\tilde{\sigma}(\cdot)$ is an online softmax operator: $m_{ij}=\max\{m_{i,j-1},\mathrm{rowmax}(S_{ij})\},~{}\widetilde{P}_{ij}=\exp(S_{ij}-m_{ij})$. Finally, the output is computed as $O_{i}=\mathrm{diag}(l_{i,n_{q}})^{-1}O_{i,n_{q}}$.
@@ 2.2 Quantization
$(\delta_{A},\hat{A})=\psi(A),~{}~{}(\delta_{B},\hat{B})=\psi(B),~{}~{}C=\psi^{-1}_{\delta_{A}\delta_{B}}(\hat{A}\hat{B})$
$\psi$ is a quantizer which converts a high-precision (e.g., FP32 or FP16) matrix $A$ to a low-precision format $\hat{A}$ (e.g., INT4 or FP8) with a scale $\delta_{A}$, and $\psi^{-1}$ is a dequantizer to convert back to high-precision. We should have $\psi^{-1}_{\delta_{A}}(\hat{A})\approx A$. The actual matrix multiplication $\hat{A}\hat{B}$ is carried out with low precision. In modern GPUs, low-precision matrix multiplication is usually multiple times faster than higher-precision ones. Quantizers differ in numerical format and granularity, e.g., how many elements (“quantization group”) share a common scale factor. For example, an INT4, per-tensor quantizer first computes the scale as the maximum absolute value of the entire tensor, scales the elements to the maximum representable range of INT4 [-7, +7], and then casts to INT4 with rounding: $\hat{A}=\lceil A/\delta_{A}\rfloor,\delta_{A}=\max(|{A}|)/7$. The dequantization process is an element-wise scaling. For example, for per-tensor dequantization, $\psi_{\delta_{A}\delta_{B}}^{-1}(\hat{A}\hat{B})=\hat{A}\hat{B}\times\delta_{A}\delta_{B}$.
@@ 2.3 SageAttention
Based on the block tiling in FlashAttention (Dao et al., 2022), SageAttention (Zhang et al., 2025) quantizes $Q,K$ to INT8 in a per-block granularity, i.e., each $Q_{i},K_{i}$ has a separate scalar scale: $\delta_{Q_{i}}=\max(|{Q_{i}|})/127,\delta_{K_{j}}=\max(|{K_{j}|})/127$. In this way, the product $S_{ij}$ in Eq. (1) can be approximated as $S_{ij}\approx\hat{Q}_{i}\hat{K}_{j}^{\top}\times(\delta_{Q_{i}}\delta_{K_{j}}/\sqrt{d})$. To maintain the accuracy, SageAttention proposes a preprocessing technique to by subtracting the token-wise mean from $K$. Additionally, SageAttention keeps $\widetilde{P}_{ij}$ and $V_{j}$ in FP16, but utilizes an FP16 accumulator (rather than FP32) for computing the product $\widetilde{P}_{ij}V_{j}$. Reducing the accumulator precision can accelerate the matrix multiplication (MM) on RTX4090 GPU. However, other GPUs, such as L20, L40, or H100, do not exhibit this behavior, as shown in Table 1.
@@ 3.1 Smooth Q
First, we discuss how to accurately compute $QK^{\top}$ with INT4. The numerical range of INT4 is notably restrictive. This affects quantization due to the presence of outliers (Lin et al., 2024). Given the INT4 range [-7, +7], any element will be quantized to zero if it is more than 14 times (0.5 vs 7) smaller than the largest element in the group. Since outliers are much larger than other elements, it is quite likely that most non-outlier elements are quantized to zero, resulting in significant accuracy degradation. Therefore, to keep the quantization accurate, we need to keep the largest element small, making the magnitude of elements as uniform as possible. Such technique is called smoothing.
Here, we propose a smoothing technique inspired by SageAttention (Zhang et al., 2025). SageAttention observed that $Q,K$ for all tokens are actually highly similar, with only small variations between different tokens (see Fig. 3). We propose to smooth $K$ as SageAttention does and further smooth $Q$ by subtracting a common mean of each block:
@@
where $\bar{q}_{i}=\mathrm{mean}(Q_{i}),\bar{k}=\mathrm{mean}(K)$ are $1\times D$ vectors, the mean is conducted along the token axis, and $\bar{q}_{i},\bar{k}$ are broadcasted to tokens in a block and a tensor for subtraction, respectively.
@@
(1) preprocessing: smooth $Q,K$ according to Eq. (2), apply quantization $(\delta_{Q_{i}},\hat{Q}_{i})=\psi_{Q}(\gamma(Q_{i})),(\delta_{K_{j}},\hat{K}_{j})=\psi_{K}(\gamma(K_{j}))$, and compute $\Delta{S_{ij}}=\bar{q}_{i}\gamma(K_{j})^{\top}$. The smoothing, quantization, and the GEMV (general matrix-vector multiplication) for computing $\Delta S$ can be fused into a single kernel, which scans the off-chip $Q$ and $K$ only once.
(2) attention: execute the low-precision GEMM, dequantize, and add back the vector $\Delta S$: $S_{ij}=\psi^{-1}_{\delta_{Q_{i}}\delta_{K_{j}}}(\hat{Q}_{i}\hat{K}_{j}^{\top})+\Delta S_{ij}$. These operations are all done on chip, and the dequantization and vector addition only adds a marginal overhead compared to the expensive mma operation for MM.
Importantly, $\gamma(Q_{i}),\gamma(K_{j})$ are quantized rather than $Q_{i},K_{j}$. Since the smoothed matrices are much smaller in magnitude and contain less outliers, the quantization accuracy can be significantly improved.
Remark. Classical techniques to improve the activation-weight MM, such as per-channel quantization, or SmoothQuant (Xiao et al., 2023) are not applicable here for the query-key MM in attention. Per-channel quantization is impracticable for $Q,K$ because quantization must be conducted along the outer axis (token dimension) of $QK^{\top}$ (Xiao et al., 2023). On the other hand, both $Q$ and $K$ have significant outliers, so trading the quantization accuracy between them with SmoothQuant cannot work effectively, as shown in Sec. 4.1. Here, we utilize the unique token-similarity pattern in attention to derive an accurate dedicated quantization method for $Q$ and $K$. The previous work SageAttention only smooths $K$, so it is less accurate than our method.
Empirical results. Fig. 19 in Appendix A.7 shows an example from CogvideoX of the distribution of $\hat{Q}$ with and without smoothing $Q$. We can find that with smoothing $Q$, the range of INT4 is utilized more uniformly and fully. Table 5 presents end-to-end metrics for different quantization methods with and without smoothing Q+K on Llama3.1 (Dubey et al., 2024) and CogvideoX (2b) (Yang et al., 2025). The results demonstrate that smoothing Q+K offers significant accuracy benefits. Also, Table 4 and 17 show that the order of effectiveness is smoothing Q+K $>$ smoothing Q $>$ smoothing K $>$ other baselines.

### v7

@@ Abstract
Although quantization for linear layers has been widely used, its application to accelerate the attention process remains limited. To further enhance the efficiency of attention computation compared to SageAttention while maintaining precision, we propose SageAttention2, which utilizes significantly faster 4-bit matrix multiplication (Matmul) alongside additional precision-enhancing techniques. First, we propose to quantize matrices $(Q,K)$ to INT4 in a hardware-friendly thread-level granularity and quantize matrices $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$, enhancing the accuracy of INT4 $QK^{\top}$. Third, we propose a two-level accumulation strategy for $\widetilde{P}V$ to enhance the accuracy of FP8 $\widetilde{P}V$. The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 4.5x. Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, while delivering much higher accuracy. Comprehensive experiments confirm that our approach incurs negligible end-to-end metrics loss across diverse models, including those for language, image, and video generation. The code is available at https://github.com/thu-ml/SageAttention.
@@ 1 Introduction
Due to the quadratic time complexity of attention, its efficient implementation becomes crucial as sequence lengths increase in real-world applications (Jiang et al., 2024; Zhang et al., 2025b). Several strategies have been developed to mitigate the computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2021; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., 2022; Xiao et al., 2024b, a; Chen et al., 2024; Jiang et al., 2024; Venkataramanan et al., 2024; Gao et al., 2024; Fu et al., 2024; Zhang et al., 2025e; Xi et al., 2025; Yang et al., 2025a; Zhang et al., 2025f, h, i) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks. The widely used attention methods optimize attention by exploiting hardware capacities, such as FlashAttention V1, V2, V3 (Dao et al., 2022; Dao, 2024; Shah et al., 2024), xformers (Lefaudeux et al., 2022), and SageAttention (Zhang et al., 2025c, d, g, a). These works do not omit computations across the entire sequence and achieve impressive speed and accuracy performance across various tasks.
@@
Motivation. For the two matrix multiplication (Matmul) operations in attention: $QK^{\top}$ and $\widetilde{P}V$, SageAttention accelerates them by quantizing the $QK^{\top}$ to INT8 and uses FP16 Matmul with FP16 accumulators for $\widetilde{P}V$. Moreover, to maintain attention accuracy, SageAttention proposes smoothing $K$ by eliminating its channel-wise outliers. SageAttention achieves 2 $\times$ speedup compared with FlashAttention2 and is the first quantized attention that incurs a negligible end-to-end metrics loss across language, image, and video generation models. However, SageAttention has two weaknesses. (W1) INT8 Matmul achieves only half the speed of INT4. (W2) FP16 Matmul with FP16 accumulators provides a speedup only on RTX 4090 and RTX 3090 GPUs. To leverage the faster INT4 tensor cores for $QK^{\top}$ and using a method that can accelerate $\widetilde{P}V$ on a broader range of GPUs, we propose to quantize $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
@@
Challenges. Quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8 presents significant challenges. For example, when only per-tensor quantizing $Q,K$ to INT4, the text-to-video model CogvideoX will generate a completely blurry video, and Llama3 only achieves a random-guessing-level accuracy of 25% on MMLU. After an in-depth investigation, we identified three primary challenges: (C1) The numerical range of INT4 is quite limited compared to FP16 or INT8, leading to significant quantization errors when $Q$ and $K$ have some abnormal values. (C2) We discover that the FP32 accumulator designed for FP8 matrix multiplication in the tensor core (mma.f32.f8.f8.f32) is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. This will lead to an accuracy loss of $\widetilde{P}V$.
@@
Our approach. To address (C1), we first discover that the per-block quantization of $Q,K$ in SageAttention does not achieve sufficient accuracy for INT4 quantization. Simultaneously, to avoid the extra latency caused by per-token dequantization, where each GPU thread corresponds to multiple quantization scales, we propose an adequately precise quantization method that incurs no additional dequantization overhead. Specifically, we introduce a per-thread quantization approach based on the mapping between the GPU threads and the memory layout of matrices as dictated by the PTX mma instructions. This method groups tokens corresponding to the same thread for quantization and dequantization, ensuring that each thread is associated with a single quantization scale. This approach achieves much better accuracy performance than per-block quantization with no additional dequantization overhead. Second, for the significant channel-wise outliers in matrices $Q$ and $K$, we adopt smoothing $K$ in SageAttention and further propose an effective method to remove these outliers in $Q$. Specifically, we propose subtracting the average value of the channel dimension of $Q$, referred to as $\overrightarrow{Q}_{m}$. Then, we add $\overrightarrow{Q}_{m}K$ after the $QK^{\top}$ Matmul to ensure the correctness of the attention computation. To address (C2), the accuracy loss caused by the 22-bit accumulator for FP8 Matmul of $\widetilde{P}V$, we propose a two-level accumulation strategy that uses an FP32 buffer to accumulate the values from the 22-bit accumulator after each block Matmul of $\widetilde{P}V$. This method confines the errors to the block range. Additionally, we design an optional technique to enhance the accuracy of the 22-bit accumulator. Specifically, we could smooth $V$ by subtracting the average value of its channel dimension and adding the subtracted item to the attention output to maintain the correctness.
@@
Performance. We offer a high-performance implementation of SageAttention2 on RTX4090 and L20 GPUs. This implementation achieves a peak performance of 481 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3x and 4.5x, respectively. To support NVIDIA Hopper GPUs, which lack native INT4 tensor core support, we additionally provide SageAttention2-8b, a variant that quantizes $Q,K$ to INT8. FlashAttention3, in contrast, is tailored to and only compatible with the Hopper architecture. Moreover, SageAttention2-8b matches the speed of FlashAttention3(fp8) on Hopper GPUs, while delivering much better accuracy. For example, on popular video generation models, our method does not compromise end-to-end accuracy, whereas FlashAttention3(fp8) brings noticeable degradation, as visualized in Fig. 7 and 9. We extensively evaluate the end-to-end metrics of state-of-the-art text, image, and video generation models. SageAttention2 can accelerate models in a plug-and-play way with negligible loss in end-to-end metrics.
@@ 2.1 FlashAttention
The attention computation can be formulated as: $S=QK^{\top}/\sqrt{d},~P=\sigma(S),~O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$. The matrices $Q$, $K$, and $V$ each have dimensionality $N\times d$, and $S$, $P$ are $N\times N$. $d$ is typically small, e.g., 64 or 128, and $N$ can be thousands or millions. The time complexity of attention is $O(N^{2})$, primarily due to two matrix multiplications ($QK^{\top}$ and $PV$), both with complexities of $O(N^{2}d)$. FlashAttention (Dao, 2024) is a GPU-friendly attention implementation, which tiles $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\}_{i=1}^{n_{q}},\{K_{i}\}_{i=q}^{n_{k}},\{V_{i}\}_{i=1}^{n_{v}}$ with block sizes of $b_{q}$, $b_{k}$, $b_{v}$ tokens, respectively, where $n_{q},n_{k},n_{v}$ are the number of tiles, and $b_{k}$ = $b_{v}$. FlashAttention computes the output matrix $O$ in parallel in tiles. Each streaming multiprocessor (SM) computes a block $O_{i}$ (corresponds to a $Q_{i}$) by iteratively loads $K_{j},V_{j}$ for each $j$, and update the output with online softmax (Milakov & Gimelshein, 2018):
@@
[…]
@@ 2.2 Quantization
[…]
$\psi$ is a quantizer which converts a high-precision (e.g., FP32 or FP16) matrix $A$ to a low-precision format $\hat{A}$ (e.g., INT4 or FP8) with a scale $\delta_{A}$, and $\psi^{-1}$ is a dequantizer to convert back to high-precision. We should have $\psi^{-1}_{\delta_{A}}(\hat{A})\approx A$. The actual matrix multiplication $\hat{A}\hat{B}$ is performed with low precision. In modern GPUs, low-precision matrix multiplication is usually multiple times faster than higher-precision ones. Quantizers differ in numerical format and granularity, e.g., how many elements (“quantization group”) share a common scale factor. For example, an INT4, per-tensor quantizer first computes the scale as the maximum absolute value of the entire tensor, scales the elements to the maximum representable range of INT4 [-7, +7], and then casts to INT4 with rounding: $\hat{A}=\lceil A/\delta_{A}\rfloor,\delta_{A}=\max(|{A}|)/7$. The dequantization process is an element-wise scaling. For example, for per-tensor dequantization, $\psi_{\delta_{A}\delta_{B}}^{-1}(\hat{A}\hat{B})=\hat{A}\hat{B}\times\delta_{A}\delta_{B}$.
@@ 2.3 SageAttention
Based on the block tiling in FlashAttention (Dao et al., 2022), SageAttention (Zhang et al., 2025c) quantizes $Q,K$ to INT8 in a per-block granularity, that is, each $Q_{i},K_{i}$ has a separate scalar scale: $\delta_{Q_{i}}=\max(|{Q_{i}|})/127,\delta_{K_{j}}=\max(|{K_{j}|})/127$. In this way, the product $S_{ij}$ in Eq. (1) can be approximated as $S_{ij}\approx\hat{Q}_{i}\hat{K}_{j}^{\top}\times(\delta_{Q_{i}}\delta_{K_{j}}/\sqrt{d})$. To maintain accuracy, SageAttention proposes a preprocessing technique by subtracting the token-wise mean from $K$. Additionally, SageAttention keeps $\widetilde{P}_{ij}$ and $V_{j}$ in FP16, but utilizes an FP16 accumulator (rather than FP32) for computing the product $\widetilde{P}_{ij}V_{j}$. Reducing accumulator precision can accelerate matrix multiplication (MM) on the RTX4090 GPU. However, other GPUs, such as L20, L40, or H100, do not exhibit this behavior, as shown in Table 1.
@@ 3.1 Smooth $Q$
First, we discuss how to accurately compute $QK^{\top}$ with INT4. The numerical range of INT4 is notably restrictive. This affects quantization due to the presence of outliers (Lin et al., 2025). Given the INT4 range [-7, +7], any element will be quantized to zero if it is more than 14 times (0.5 vs 7) smaller than the largest element in the group. Since outliers are much larger than other elements, it is likely that many non-outlier elements are quantized to zero, leading to substantial accuracy degradation. Therefore, to keep the quantization accurate, we need to keep the largest element small, making the magnitude of elements as uniform as possible. Such technique is called smoothing.
Here, we propose a smoothing technique inspired by SageAttention (Zhang et al., 2025c). SageAttention observed that $Q,K$ for all tokens are actually highly similar, with only small variations between different tokens (Fig. 2 shows the heatmap distributions of the $Q$, $K$, and $V$ randomly sampled from Llama3.1 (Dubey et al., 2024) and CogvideoX (Yang et al., 2025b)). We propose to smooth $K$ as SageAttention does and further smooth $Q$ by subtracting a common mean of each block:
@@
where $\bar{q}_{i}=\mathrm{mean}(Q_{i}),\bar{k}=\mathrm{mean}(K)$ are $1\times D$ vectors, the mean is conducted along the token axis, and $\bar{q}_{i},\bar{k}$ are broadcasted to tokens in a block and a tensor for subtraction.
@@
(1) preprocessing: smooth $Q,K$ according to Eq. (2), apply quantization $(\delta_{Q_{i}},\hat{Q}_{i})=\psi_{Q}(\gamma(Q_{i})),(\delta_{K_{j}},\hat{K}_{j})=\psi_{K}(\gamma(K_{j}))$, and compute $\Delta{S_{ij}}=\bar{q}_{i}\gamma(K_{j})^{\top}$. Smoothing, quantization, and GEMV (general matrix-vector multiplication) for computing $\Delta S$ can be fused into a single kernel, which reads the off-chip $Q$ and $K$ only once.
(2) attention: execute the low-precision GEMM, dequantize, and add back the vector $\Delta S$: $S_{ij}=\psi^{-1}_{\delta_{Q_{i}}\delta_{K_{j}}}(\hat{Q}_{i}\hat{K}_{j}^{\top})+\Delta S_{ij}$. These operations are all done on chip, and the dequantization and vector addition only add a marginal overhead compared to the expensive mma operation for MM.
Importantly, $\gamma(Q_{i}),\gamma(K_{j})$ are quantized rather than $Q_{i},K_{j}$. Since the smoothed matrices are much smaller in magnitude and contain fewer outliers, the quantization accuracy can be significantly improved. A theoretical analysis of the benefit of smoothing is included in Appendix A.5.
Remark. Classical techniques to improve the activation-weight MM, such as per-channel quantization, or SmoothQuant (Xiao et al., 2023) are not applicable here for the query-key MM in attention. Per-channel quantization cannot be applied to $Q,K$ because the quantization must be conducted along the outer axis (token dimension) of $QK^{\top}$. On the other hand, both $Q$ and $K$ have significant outliers, so trading the quantization accuracy between them with SmoothQuant cannot work effectively, as shown in Sec. 4. Here, we utilize the unique token similarity pattern in attention to derive a dedicated quantization method for $Q$ and $K$. The previous work SageAttention only smooths $K$, so it is less accurate than our method.
Empirical results. Fig. 20 in Appendix A.9 shows an example from CogvideoX of the distribution of $\hat{Q}$ with and without smoothing $Q$. We can find that with smoothing $Q$, the range of INT4 is utilized more uniformly and fully. Table 5 presents end-to-end metrics for different quantization methods with and without smoothing Q+K on Llama3.1 and CogvideoX (2b). The results demonstrate that smoothing Q+K offers significant accuracy benefits. Also, Table 4 and 17 show that the order of effectiveness is smoothing Q+K >> smoothing Q >> smoothing K >> other baselines.